The sum of ages of father and a son is 45 years.5 years ago the product of their ages was 4 times the fathers age at that time.Find the present ages of father and son?
Answer by stanbon(75887)

The sum of ages of father and a son is 45 years.5 years ago the product of their ages was 4 times the fathers age at that time.Find the present ages of father and son?
---
Equations:
f + s = 45
(f-5)(s-5)=4(f-5)
----
Substitute for "f" and solve for "s":
f = 45-s
(45-s-5)(s-5) = 4(45-s-5)
---
Divide both sides by 45-s-5 to get:
s-5 = 4
s = 9 (son's age now)
----
Solve for "f":
f+9 = 45
f = 36 (father's age now)
============================
Cheers,
Stan H.